# Edge-Cloud Routing for Text-to-Image Model with Token-Level Multi-Metric Prediction

### v1

@@ Abstract
Large text-to-image models demonstrate impressive generation capabilities; however, their substantial size necessitates expensive cloud servers for deployment. Conversely, light-weight models can be deployed on edge devices at lower cost but often with inferior generation quality for complex user prompts. To strike a balance between performance and cost, we propose a routing framework, called RouteT2I, which dynamically selects either the large cloud model or the light-weight edge model for each user prompt. Since generated image quality is challenging to measure directly, RouteT2I establishes multi-dimensional quality metrics, particularly, by evaluating the similarity between the generated images and both positive and negative texts that describe each specific quality metric. RouteT2I then predicts the expected quality of the generated images by identifying key tokens in the prompt and comparing their impact on the quality. RouteT2I further introduces the Pareto relative superiority to compare the multi-metric quality of the generated images. Based on this comparison and predefined cost constraints, RouteT2I allocates prompts to either the edge or the cloud. Evaluation reveals that RouteT2I significantly reduces the number of requesting large cloud model while maintaining high-quality image generation.
@@ 1 Introduction
Text-to-Image Model #Parameter Pricing (\$/M)
DALL·E [24] 12 B -
DALL·E 2 [23] 3.5 B 20 K
DALL·E 3 [4] - 80 K
Imagen [3] - 50 K
@@
Parti [32] 20 B -
Playground v3 [16] 24 B -
FLUX 1.1 [1] 12 B 40 K
Table 1: Overview of popular text-to-image models, including their parameters and the cost of generating millions of images.
Nowadays, text-to-image (T2I) models like Imagen [3], Stable Diffusion [25], and DALL·E [24] have achieved significant success in generating diverse, high-quality images given user prompts. However, the impressive generation quality comes with large model and high cost. For example, Stable Diffusion 3.5 [2] has 8 billion parameters. Such a large model necessitates reliance on cloud servers, leading to high serving cost. As shown in Tab. 1, it is particularly costly in commercial scenarios with millions of requests [^1] https://artificialanalysis.ai/text-to-image.
@@
The high cost of cloud-based image generation drives the trend of deploying T2I models on edge devices. Techniques such as quantization [13, 34, 30], structural pruning [14, 5], and reducing denoising steps [19, 17, 27] are used to minimize model sizes and speed up inference, showing that edge deployment is feasible. Compared to cloud models, edge models leverage users’ local computing devices to provide instant services anytime and anywhere without cloud serving cost and communication overhead. However, the lightweight edge T2I models often have lower generation quality compared to the cloud large models.
@@
Although cloud models often offer superior quality, not all user prompts need these large models. For simple user prompts, lightweight edge models can produce comparable or even better results. To ensure high-image quality generation while limiting requests to cloud models to reduce cost, it is necessary to introduce a routing mechanism such that suitable models are selected for different user prompts based on their complexity. Intuitively, as shown in Fig. 1, a routing framework for T2I models can route only hard prompts to the cloud, while handling easy ones on the edge.
@@
Previous routing methods are primarily designed for classification models [11] and large language models (LLMs) [18, 20, 7, 9], presenting significant challenges when applied to T2I models. One major challenge lies in measuring the quality of generated images. Unlike text, image quality does not have unified metrics and is subject to many influencing factors, including color, clarity, and the unique distortions typical of generated images [12]. This complicates the objective and comprehensive evaluation of generated image quality. Furthermore, the output space of T2I models is substantially larger than the input text space, and a single text description can correspond to numerous images, hindering the effective prediction of image quality based on text prompts.
Figure 1: Edge-cloud routing for text-to-image generation models.
@@
To address these challenges, we propose a new edge-cloud T2I routing framework RouteT2I, deploying a routing model and a small T2I generation model on the edge, and a large T2I generation model on the cloud. The design goal is to optimize the overall quality of generated images at a limited cost, by effectively routing user prompts to the edge or the cloud. To evaluate the quality gap between generated images, we first propose a multi-metric quality measure, including metrics for both traditional photos and those unique to generated images. Each metric assesses image quality by measuring the similarity between the image and a pair of positive and negative prompts that describe the quality metric. We then define Pareto relative superiority (PRS) to quantify quality differences between generated images. RouteT2I predicts PRS based on user prompts, inspired by the cross-attention operation between text and image during generation. In particular, we design a novel Mixture-of-Experts (MoE) [28] network, dual-gate token selection MoE, where a user prompt is treated as a sequence of tokens, with experts aligned to quality metrics. A token selection gate allows experts to proactively choose tokens that markedly impact these metrics, thereby focusing on key tokens. A dual-gate MoE leverages both positive and negative gates to evaluate dominant influences when tokens have opposing effects simultaneously. Furthermore, multiple heads are introduced to predict multiple quality metrics. Based on the multi-metric quality prediction, RouteT2I determines quality differences between images generated on the edge over those in the cloud. Only user prompts that show substantial quality improvement when generated in the cloud are routed there, while others remain at the edge.
We summarize the key contributions as follows:
  • We, for the first time, consider and formulate the problem of text-to-image model routing between cloud and edge.
@@
  • We propose a contrastive multi-metric quality measure method for generated images and further design a routing model architecture with dual gate token selection MoE coupled with a routing strategy for user prompts.
@@
  • We evaluate the proposed design using the public COCO2014 dataset [15] and 10 pairs of edge-cloud generation models. Key results include a relative quality improvement of 83.97% at a routing rate of 50%, and the cloud request reduction of 70.24% at the quality target of 50%.
@@ 2 Related Work
Previous model routing work primarily focused on classification models and large language models (LLMs). Depending on whether routing decision occurs before or after weak model execution, model routing can be categorized into predictive and non-predictive routing.
Non-predictive routing takes the output of a weaker model to decide whether to route to the next, more powerful model. LAECIPS [10] evaluates the confidence of predictions made by the edge semantic segmentation model to determine whether assistance from the cloud’s large visual model is necessary. For classification models, Hybrid [11] employed a routing model to partition the data domain between edge and cloud models, routing cases with incorrect edge outputs to the cloud. In the context of LLM, LLM Cascade [33] routed based on the consistency of responses from weaker LLMs, while FrugalGPT [6] considered current quality and historic total costs in the cascading LLM to decide on routing. Tabi [31] not only routed according to the confidence of responses from weaker LLMs but also enhanced output quality by aggregating historic outputs.
Predictive routing selects the appropriate model before running the initial, weaker model. Previous studies focused on LLM routing, with the difference in routing strategies. RouteLLM [20] predicted the outcomes of quality comparisons between LLM outputs and routed based on prediction confidence. Hybrid LLM [7] relaxed the comparison criteria, allowing the weak model to succeed if the quality gap is within a threshold, thus saving costs with some quality compromise. Recent research [29] advances the confidence routing strategy by identifying out-of-distribution data through confidence thresholds. ZOOTER [18] introduced a new routing strategy by predicting the normalized quality of candidate model outputs, making routing decision based on relative quality, with routing models distilled from existing quality scoring models.
@@ 3 Problem Formulation
In text-to-image generation scenario, the cloud normally hosts a large T2I model, denoted as $\mathcal{M}_{c}$, while a resource-constraint edge device, such as a smartphone, often deploys a lightweight T2I model, denoted as $\mathcal{M}_{e}$. For example, Google has released Imagen [26, 3] for cloud deployment and MobileDiffusion [35] for edge deployment. From the perspective of a user, the large model generally offers superior generation quality, but incurs edge-cloud communication cost and cloud serving fee, which together are denoted as $F_{c}$. In contrast, the edge model can leverage the user’s local device, thereby eliminating these additional costs.
In practice, user requests involve not only complex ones but also simple ones. For some requests, particularly simple text prompts, the performance gap between the large model and the light-weight model is negligible, and sometimes, the lightweight model even performs better. In such cases, the purely large model serving not only fails to improve performance but also incurs high cost. Therefore, it is necessary to design a routing mechanism to select the most suitable model based on the complexity of the prompt. Intuitively, the desired routing mechanism should direct simpler prompts to the cost-effective edge model and more complex prompts to the high-quality cloud model, thus maximizing generation quality while minimizing cost.
@@
$\max\ &R(\mathcal{X})\mathcal{Q}(\mathcal{I}_{c})+(1-R(\mathcal{X}))\mathcal{Q}(\mathcal{I}_{e})\\ \text{s.t.}\ &\mathbb{P}\{R(\mathcal{X})=1\}\cdot F_{c}\leq\tau_{fee}\\ &\mathbb{P}\{R(\mathcal{X})=1\}\cdot D_{\mathcal{M}_{c}}+\mathbb{P}\{R(\mathcal{X})=0\}\cdot D_{\mathcal{M}_{e}}\\ &+D_{R}\leq\tau_{time},\end{aligned}$ (1)
where $D_{\mathcal{M}_{c}}$, $D_{\mathcal{M}_{e}}$, and $D_{R}$ denote the latency of cloud large model serving plus edge-cloud communication, the latency of the edge light-weight model serving, and the latency of the routing model execution. Considering the constraints are linear, the optimization objective can be simplified to imposing an upper bound $\rho_{r}$ on the routing rate to the cloud large model as
$\max\ &R(\mathcal{X})\mathcal{Q}(\mathcal{I}_{c})+(1-R(\mathcal{X}))\mathcal{Q}(\mathcal{I}_{e})\\ \text{s.t.}\ &\mathbb{P}\{R(\mathcal{X})=1\}\leq\rho_{r}.\end{aligned}$ (2)
Due to the lack of a universal standard for evaluating the quality of generated images, a multi-dimensional quality scoring function $\mathcal{Q}:\mathcal{I}\rightarrow[0,1]^{N}$ is required for a more comprehensive comparison. As a result, different from existing routing frameworks for typical classification and text generation tasks, the T2I routing framework’s optimization objective is a new multi-objective problem. On the other hand, from the perspective of input-to-output mapping complexity, the considered T2I task presents a unique challenge, because the text input space is significantly smaller than the image output space. In contrast, the classification task has a more constrained output space, while the text generation task involves comparable input and output spaces, rendering previous routing designs inapplicable for T2I tasks. By examining the interaction between text and images during the T2I generation process, we conduct a detailed analysis of how different prompt tokens influence the generated image and its multi-dimensional quality metrics. In what follow, we define the contrastive multi-metric quality measure in Sec. 4 and introduce the T2I routing framework in Sec. 5.
4 Contrastive Multi-Metric Image Generation Quality Measure
The connection between text and image enables the depiction of an image’s properties through corresponding textual descriptions. We thus propose a quality metric by evaluating the alignment between the text prompt that describe this metric and the image, typically, the cosine similarity between text features and image features using CLIP [22]. To enhance the accuracy and stability of image quality measure, we introduce positive and negative text prompts that describe opposing levels of the quality metric for contrast. The contrastive quality metric of image $I$ is defined as
@@ 4 Contrastive Multi-Metric Image Generation Quality Measure
Considering that the quality of real photos relies on many factors, such as definition and color, we introduce multi-dimensional metrics to comprehensively measure the quality of generated images, offering stability against noise and uncertainty. In addition to the above metrics for real photos, we also incorporate unique metrics specifically for generated images, such as realism and object integrity, to establish a comprehensive set of multiple metrics. For each original quality metric $i$ out of all the $N$ metrics, we design a pair of negative and positive prompts $m_{i}=(m_{i}^{+},m_{i}^{-})$ that describe $i$ in text. The $N$-dimensional contrastive quality metrics for an image $I$ can be expressed as
@@
where $m_{i}$ is an instantiation of $m$ in Eq. 3. For example, for the quality metric of definition, we use the positive prompt “High definition photo” and the negative prompt “Low definition photo”; and for the the quality metric of object integrity, we take the positive prompt “Object completion” and the negative prompt “Object twisting”.
In the multi-objective optimization problem setting for routing with multi-metric quality, the initial goal is to find a Pareto optimal generated image that excels in all the metrics. However, this is challenging in practice due to the potential nonexistence of such an image generated from the cloud or edge model. We thus relax the constraint of Pareto optimality by allowing suboptimal performance in some metrics if the image significantly outperforms in others when selecting higher-quality images. We define Pareto relative superiority (PRS) to quantify the quality advantage of the images $\mathcal{I}_{e}$ generated by the edge model over the images $\mathcal{I}_{c}$ generated by the cloud model. In particular, we first normalize the quality distance for the metric $i$ between an edge-generated image $I_{e}\in\mathcal{I}_{e}$ and a cloud-generated image $I_{c}\in\mathcal{I}_{c}$ as
@@
where $w_{i}$ denotes the importance weight of the metric $i$, and $\sum_{i}w_{i}=1$. We can evaluate the relative quality of $I_{e}$ compared to $I_{c}$ by examining how much Pareto relative superiority deviates from 0.5. This evaluation helps to route a user prompt to the model that generates an image with higher overall quality.

### v2

# Adaptive Routing of Text-to-Image Generation Requests Between Large Cloud Model and Light-Weight Edge Model

@@ Abstract
Large text-to-image models demonstrate impressive generation capabilities; however, their substantial size necessitates expensive cloud servers for deployment. Conversely, light-weight models can be deployed on edge devices at lower cost but often with inferior generation quality for complex user prompts. To strike a balance between performance and cost, we propose a routing framework, called RouteT2I, which dynamically selects either the large cloud model or the light-weight edge model for each user prompt. Since generated image quality is challenging to measure and compare directly, RouteT2I establishes multi-dimensional quality metrics, particularly, by evaluating the similarity between the generated images and both positive and negative texts that describe each specific quality metric. RouteT2I then predicts the expected quality of the generated images by identifying key tokens in the prompt and comparing their impact on the quality. RouteT2I further introduces the Pareto relative superiority to compare the multi-metric quality of the generated images. Based on this comparison and predefined cost constraints, RouteT2I allocates prompts to either the edge or the cloud. Evaluation reveals that RouteT2I significantly reduces the number of requesting large cloud model while maintaining high-quality image generation.
@@ 1 Introduction
Nowadays, text-to-image (T2I) models like Imagen [3], Stable Diffusion [28], and DALL·E [27] have achieved significant success in generating diverse, high-quality images given user prompts. However, the impressive generation quality comes with large model and high cost. For example, Stable Diffusion 3.5 [1] has 8 billion parameters. Such a large model necessitates reliance on cloud servers, leading to high serving cost. As shown in Tab. 1, it is particularly costly in commercial scenarios with millions of requests.
Text-to-Image Model #Param Pricing (\$/M)
@@
Table 1: For the Stable Diffusion models, newer versions come with more parameters and higher API costs.
@@
The high cost of cloud-based image generation drives the trend of deploying T2I models on edge devices. Techniques such as quantization [16, 43, 34], structural pruning [17, 4], and reducing denoising steps [22, 20, 30] are used to minimize model sizes and speed up inference, showing that edge deployment is feasible. Compared to cloud models, edge models leverage users’ local computing devices to provide instant services anytime and anywhere without cloud serving cost and communication overhead. However, the light-weight edge T2I models often have lower generation quality compared to the cloud large models.
@@
Although cloud models often offer superior quality, not all user prompts need these large models. For some prompts, light-weight edge models can produce comparable or even better results. To ensure high-image quality generation while limiting requests to cloud models to reduce cost, it is necessary to introduce a routing mechanism such that suitable models are selected for different user prompts based on their complexity. Intuitively, as shown in Fig. 1, a routing framework for T2I models routes only challenging prompts to the cloud, while handling easy ones on the edge.
@@
Previous routing methods are primarily designed for large language models (LLMs) [21, 23, 8, 11], presenting significant challenges when applied to T2I models. One major challenge lies in comparing the quality of generated images. Unlike text, image quality does not have unified metrics and is subject to many influencing factors, including color, clarity, and the unique distortions typical of generated images [15]. This makes it hard to compare image quality to guide routing decisions. Furthermore, the output space of T2I models is substantially larger than the input text space, and a single prompt can correspond to numerous images, hindering the prediction of image quality before generation.
Figure 1: Routing for text-to-image generation requests.
@@
To address these challenges, we propose an edge-cloud T2I routing framework RouteT2I, deploying a routing model and a small T2I model on the edge, and a large T2I model on the cloud. Our goal is to optimize the overall quality of generated images at a limited cost, by effectively routing user prompts to the edge or the cloud. To evaluate the quality gap between generated images, we first propose a multi-metric quality measure, including metrics for both real photos and those unique to generated images. Each metric assesses image quality by measuring the similarity between the image and a pair of positive and negative prompts that describe the quality metric. We then define Pareto relative superiority (PRS) to quantify quality differences between images. RouteT2I predicts PRS based on user prompts, inspired by the cross-attention operation between text and image during generation. In particular, inspired by mixture-of-experts (MoE) selective focus on experts or tokens in LLMs [32, 45, 2], we design a novel dual-gate token selection MoE, where a user prompt is treated as a sequence of tokens, with experts aligned to quality metrics. A token selection gate allows experts to proactively choose tokens that markedly impact these metrics, thereby focusing on key tokens. A dual-gate MoE uses both positive and negative gates to evaluate dominant effects when tokens have opposing effects simultaneously. Furthermore, multiple heads are introduced to predict multiple quality metrics. Based on the multi-metric quality prediction, RouteT2I determines quality differences between images generated on the edge over those in the cloud. Only user prompts that show substantial quality improvement when generated in the cloud are routed there, while others remain at the edge. Our key contributions are as follows:
  • We, for the first time, study how to route text-to-image generation requests between cloud and edge models for low cost and high quality.
@@
  • Specified to multi-metric quality measures of generated images, we design a routing model architecture with dual gate token selection MoE, combined with a routing strategy for user prompts.
@@
  • We evaluate the proposed design using 18 image generation model pairs on the public COCO 2014 dataset. Key results include: (1) by routing 50% of requests to edge and cloud models, respectively, the overall quality improvement over the edge model reaches 83.97% of that achieved by fully using the cloud model; and (2) to achieve 50% of the quality improvement attained by fully using the cloud model, the number of requests routed to the cloud model is reduced by 70.24% compared to a random routing policy.
@@ 2 Related Work
Model routing selects suitable models for user requests to ensure quality while reducing high cost of invoking powerful models. Depending on whether routing decision occurs before or after weak model execution, model routing can be categorized into predictive and non-predictive routing.
Non-predictive routing takes the output of a weak model to decide whether to route to the next, powerful model. LAECIPS [12] evaluated the confidence of predictions made by the edge semantic segmentation model to decide if cloud assistance from a large visual model is needed. For classification models, Hybrid [13] used a routing model to partition the data domain between edge and cloud models, routing cases with incorrect edge outputs to the cloud. For cascading LLM, LLM Cascade [41] routed based on the consistency of responses from weaker LLMs, while FrugalGPT [6] considered current quality and past costs to determine routing. Tabi [36] not only routed based on the confidence of responses from weaker LLMs but also enhanced output quality by aggregating historic outputs.
Different from non-predictive routing, predictive routing selects model before running the weak model on resource-constrained edge device. Previous studies focused on LLM, with the difference in routing strategies. RouteLLM [23] predicted the outcomes of quality comparisons between LLM and routed based on prediction confidence. Hybrid LLM [8] relaxed comparisons, allowing the weak model to succeed if the quality gap is within a threshold, thus saving costs with some quality compromise. Recent research [33] advances the confidence routing strategy by identifying out-of-distribution data through confidence thresholds. ZOOTER [21] introduced a new strategy by predicting the normalized quality of candidate model outputs, making routing decision based on relative quality, with routing models distilled from existing quality scoring models.
Previous routing work were mainly for LLMs, whereas the predictive routing for T2I in our work remains unexplored. A key distinction lies in the image outputs of T2I, which necessitate predicting quality variations in a larger output image space from a smaller input text space (Sec. B in Appendix). On one hand, unlike many LLM tasks with definitive answers, image quality lacks fixed standards and is highly subjective. Previous single-objective LLM routing struggles to adapt to such ambiguous quality criteria. On the other hand, prior work lacked dedicated routing model designs, making it challenging to predict quality relationships within the vast output image space. This underscores the necessity of effective routing model design.
@@ 3 Problem Formulation
In text-to-image generation scenario, the cloud normally hosts a large T2I model, denoted as $\mathcal{M}_{c}$, while a resource-constraint edge device, such as a smartphone, often deploys a light-weight T2I model, denoted as $\mathcal{M}_{e}$. For example, Google has released Imagen [29, 3] for cloud deployment and MobileDiffusion [44] for edge deployment. From the perspective of a user, the large model generally offers superior generation quality, but incurs edge-cloud communication cost and cloud serving fee, which together are denoted as $F_{c}$. In contrast, the edge model can leverage the user’s local device, thereby eliminating these additional costs.
Figure 2: SD2.1 tends to match or outperform SD3 on simple prompts, while more likely falling short on complex ones. Entities to be generated are bolded.
Figure 3: Impact of the noun count in text prompts on ImageReward score [40] gap and win rate between SD3 and SD2.1.
In practice, user requests involve not only complex ones but also simple ones. One intuitive factor influencing request complexity is the number of entities that need to be generated in images, i.e. the number of nouns in text prompts. As shown in Fig. 2, 3, as noun count increases, the large model is more likely to outperform the small model in quality, with quality differences becoming more noticeable. However, while large models generally excel in overall quality, small models may perform better on specific prompts. This likelihood increases as the prompt becomes simpler. In such cases, the purely large model serving not only fails to improve performance but also incurs high cost. Therefore, it is necessary to design a routing mechanism to select the most suitable model based on the prompt’s characteristics. Intuitively, the routing mechanism should direct simpler prompts to the cost-effective edge model and more complex prompts to the high-quality cloud model, thus maximizing generation quality while minimizing cost.
@@
$\begin{aligned} \max\ &R(\mathcal{X})\mathcal{Q}(\mathcal{I}_{c})+(1-R(\mathcal{X}))\mathcal{Q}(\mathcal{I}_{e})\\ \text{s.t.}\ &\mathbb{P}\{R(\mathcal{X})=1\}\cdot F_{c}\leq\tau_{fee}\\ &\mathbb{P}\{R(\mathcal{X})=1\}\cdot D_{\mathcal{M}_{c}}+\mathbb{P}\{R(\mathcal{X})=0\}\cdot D_{\mathcal{M}_{e}}\\ &+D_{R}\leq\tau_{time},\end{aligned}$ (1)
where $D_{\mathcal{M}_{c}}$, $D_{\mathcal{M}_{e}}$, and $D_{R}$ denote the latency of cloud large model serving plus edge-cloud communication, the latency of the edge light-weight model serving, and the latency of the routing model execution. Considering the constraints are linear, the optimization objective can be approximated as imposing an upper bound $\rho_{r}$ on the routing rate to the cloud large model as
$\begin{aligned} \max\ &R(\mathcal{X})\mathcal{Q}(\mathcal{I}_{c})+(1-R(\mathcal{X}))\mathcal{Q}(\mathcal{I}_{e})\\ \text{s.t.}\ &\mathbb{P}\{R(\mathcal{X})=1\}\leq\rho_{r}.\end{aligned}$ (2)
Due to the lack of a universal standard for evaluating the quality of generated images, a multi-dimensional quality scoring function $\mathcal{Q}:\mathcal{I}\rightarrow[0,1]^{N}$ is required for a more comprehensive comparison. Thus, different from existing routing for typical classification and text generation tasks, the T2I routing framework’s optimization objective is a new multi-objective problem. In addition, to predict quality changes in the larger image output space from the smaller text input space, we conduct a detailed analysis of how different prompt tokens influence the generated images and their multi-dimensional quality metrics, by examining the interaction between text and images during the T2I generation process. In what follow, we define the multi-metric quality measures for routing objectives in Sec. 4 and introduce the T2I routing framework in Sec. 5.
4 Multi-Metric Image Generation Quality in Routing Objectives
The connection between text and images allows image attributes to be described through corresponding prompts. Thus, a metric can be measured by the relative relationship between an image and the positive/negative text pair describing that metric [35, 18]. The contrastive quality metric for image $I$ is defined as
@@ 4 Multi-Metric Image Generation Quality in Routing Objectives
Considering the ambiguity and complexity of image quality, we introduce a multi-dimensional metric to comprehensively evaluate generated image quality, providing noise resistance and stability for quality-prediction-based routing. Just as real photo quality depends on factors like definition and color, our multi-metric combines traditional measures of real photos with unique aspects of generated images, such as realism and object integrity. Thus, we extend the single-metric evaluation in Eq. 3 to an $N$-dimensional quality assessment of an image $I$, expressed as
@@
where $m_{i}$ is an instantiation of $m$ in Eq. 3. We adopt 10 common metrics inspired by prior works [14, 37, 39], with details in Appendix Sec. C. For example, for the definition metric, we use the positive prompt “High definition photo” and the negative prompt “Low definition photo”; and for the object integrity, we take the positive prompt “Object completion” and the negative prompt “Object twisting”.
In the multi-objective optimization setting for routing with multi-metric quality, the initial goal is to find a Pareto optimal generated image that excels in all the metrics. However, this is challenging in practice due to the potential nonexistence of such an image generated from the cloud or edge model. We thus relax the constraint of Pareto optimality by allowing suboptimal performance in some metrics if the image significantly outperforms in others when selecting higher-quality images. We define Pareto relative superiority (PRS) to quantify the quality advantage of the images $\mathcal{I}_{e}$ generated by the edge model over the images $\mathcal{I}_{c}$ generated by the cloud model. In particular, we first normalize the quality distance for the metric $i$ between an edge-generated image $I_{e}\in\mathcal{I}_{e}$ and a cloud-generated image $I_{c}\in\mathcal{I}_{c}$ as
@@
where $w_{i}$ denotes the weight of metric $i$, and $\sum_{i}w_{i}=1$. We can evaluate the relative quality of $I_{e}$ compared to $I_{c}$ by examining how much Pareto relative superiority deviates from 0.5. This evaluation helps to route a user prompt to the model that generates an image with higher overall quality.